# On the existence of forward self-similar solutions to the two-dimensional Navier-Stokes equations

### v1

@@ 2. The Leray equations and the Energy estimates
Furthermore, we demonstrate that $u_{\text{re}}(x,1)$ is $L^{p}$-integrable for any $p\in(1,+\infty)$. This implies that $u_{\textrm{re}}$ decays like $\frac{1}{1+|x|^{2}}$ in an average sense. The local regularity of $u_{\text{re}}(x,1)$ and $e^{\Delta}u_{0}$ is not a big issue, indeed they are smooth due to the parabolic smoothing effect, once we have some initial local integrability. However, they decay slowly at spatial infinity, and the key point is to get the global integrability to initiate the subsequent analysis. Before stating the main result in this section, we first introduce self-similar variables and the Leray equations.
@@ 2.1. Self-similar variables and the Leray equations
Our analysis will be conducted in self-similar variables:
(2.1) $y=\frac{x}{\sqrt{t}},~s=\log(t).$
We define $v(y,s)$ as
@@
(2.3) $\left\{\begin{aligned} &\frac{\partial v}{\partial s}-\Delta v-\frac{1}{2}v-\frac{1}{2}y\cdot\nabla v+v\cdot\nabla v+\nabla q=0,\\ &\textrm{ div }v=0.\end{aligned}\right$.
@@
(2.4) $\left\{\begin{aligned} &-\Delta v-\frac{1}{2}v-\frac{1}{2}y\cdot\nabla v+v\cdot\nabla v+\nabla q=0,\\ &\textrm{ div }v=0.\end{aligned}\right$.
@@
$e^{t\Delta}u_{0}(x,t)=\int_{\mathbb{R}^{2}}\frac{1}{4\pi t}e^{-\frac{|x-z|^{2}}{4t}}u_{0}(z)dz.$
Direct calculation shows that $e^{t\Delta}u_{0}$ inherits the self-similarity of $u_{0}$:
$\lambda(e^{t\Delta}u_{0})(\lambda x,\lambda^{2}t)$ $=\lambda\int_{\mathbb{R}^{2}}\frac{1}{4\pi\lambda^{2}t}e^{-\frac{|\lambda x-z|^{2}}{4\lambda^{2}t}}u_{0}(z)dz$
@@
$=\int_{\mathbb{R}^{2}}\frac{1}{4\pi t}e^{-\frac{|x-z|^{2}}{4t}}u_{0}(z)dz=e^{t\Delta}u_{0}(x,t),$
@@
$v_{0}(y)=\sqrt{t}(e^{t\Delta}u_{0})(x,t),$
then obviously we have
(2.5) $v_{0}(y)=(e^{t\Delta}u_{0})\left(\frac{x}{\sqrt{t}},1\right)=(e^{\Delta}u_{0})(y).$
@@
Our primary goal will then be to get a priori estimates for $v_{\text{re}}(y)$. It is easy to see that $v_{0}$ satisfies
@@
(2.8) $\left\{\begin{aligned} &-\Delta v_{\text{re}}-\frac{1}{2}v_{\text{re}}-\frac{1}{2}y\cdot\nabla v_{\text{re}}+v_{0}\cdot\nabla v_{\text{re}}+v_{\text{re}}\cdot\nabla v_{0}+v_{\text{re}}\cdot\nabla v_{\text{re}}+\nabla q=-v_{0}\cdot\nabla v_{0},\\ &{\rm div\thinspace}v_{\text{re}}=0.\end{aligned}\right$.
The main objective of this Section is to prove the following.
@@
Let the divergence-free, self-similar initial data $u_{0}(x)\in C^{0,\beta}_{loc}$ be given, $0<\beta\leq 1$. We denote
@@
Assume $u$ is a self-similar solution to (1.1) in the sense of Definition 1.1. Consider the corresponding solution to (2.8) in which $v_{0}=e^{\Delta}u_{0}$, then there exists a constant $C=C(A,\beta)$ depending only on $A$ and $\beta$, such that
$\int_{\mathbb{R}^{2}}|v_{\text{re}}|^{2}+|\nabla v_{\text{re}}|^{2}dx\leq C(A,\beta).$
@@
$\|\nabla q\|_{L^{2}(\mathbb{R}^{2})}\leq C(A,\beta).$
@@ 2.2. Main difficulties, key ideas and the outline for the energy estimates
In this section, we give the main idea for the a priori estimates of solutions in Theorem 2.1.
@@
yet $\|v_{\text{re}}\|_{L^{2}(\mathbb{R}^{2})}$ cannot be controlled at this stage. This situation differs from the three-dimensional case, where there is an additional term $\frac{1}{4}\int_{\mathbb{R}^{2}}|v_{\text{re}}|^{2}$ that appears in the left-hand side of (2.9) and help to close the energy estimates, provided $\|v_{0}\|_{L^{\infty}}$ is small via a suitable cut-off; see [4, 5].
@@
as $\nabla v_{0}\sim\frac{C}{1+|y|^{1+\beta}}\in L^{2}(\mathbb{R}^{2})$. This provides the initial Dirichlet estimate for $v_{\text{re}}$.
@@
The left hand side contains an additional term $\int_{\mathbb{R}^{2}}\left(\frac{1}{p}-\frac{1}{2}\right)|v_{\text{re}}|^{p}$, which enables us to close the estimates for $\|v_{\text{re}}\|_{L^{p}}$ for any $1<p<2$ eventually. The key issue is to control $\int_{\mathbb{R}^{2}}(v_{\text{re}}\cdot\nabla v_{0})\cdot|v_{\text{re}}|^{p-2}v_{\text{re}}$ and $\int_{\mathbb{R}^{2}}\nabla q\cdot|v_{\text{re}}|^{p-2}v_{\text{re}}$ as the last term on the right hand side is more manageable. To achieve this, we introduce a modified profile of $v_{0}$ to ensure it is small near the origin. Finally, once the $\|v_{\text{re}}\|_{L^{p}}$ estimate is established, the Gagliardo–Nirenberg interpolation inequality combined with (2.11) yields the desired $L^{2}$-boundedness of $v_{\text{re}}$. The higher integrability estimate of $\|v_{\text{re}}\|_{L^{p}}$ when $p>2$ will follow from regularity estimates for the Stokes equation and the bootstrapping argument.
@@ 2.3. Proof of the energy estimates
We need to do some preliminary work. We begin with the following decay estimates for the solution to the heat equation. It is more or less standard, and we leave the proof to Appendix A.
@@
Let $u_{0}$ be divergence-free on $\mathbb{R}^{2}$ and self-similar, and let $0<\beta\leq 1$, denote
@@
Then $v_{0}=e^{\Delta}u_{0}(y)$ defined in (2.5) is smooth and divergence-free.
@@
so that there exists some universal constant $C$ that
@@
where $w$ is used to make $v_{1}$ is divergence-free, so $w$ solves
${\rm div\thinspace}w=-v_{0}\cdot\nabla\eta_{R_{0}}.$
Explicitly, we can define
(2.15) $w(y)$ $=-\frac{1}{2\pi}\nabla_{y}\int_{\mathbb{R}^{2}}(\ln|y-z|)v_{0}\cdot\nabla\eta_{R_{0}}(z)dz$
$=-\frac{1}{2\pi}\int_{\mathbb{R}^{2}}\frac{y-z}{|y-z|^{2}}v_{0}\cdot\nabla\eta_{R_{0}}(z)dz.$
@@
Such a correction term $w$ is obviously not unique. Typically, it can be constructed with compact support in the annulus $B_{2R_{0}}\setminus B_{R_{0}}$ using the Bogovskii formula, given that $v_{0}\cdot\nabla\eta_{R_{0}}$ is smooth and supported in that region (cf. [13, Theorem III.3.3]). The representation in (2.15), however, may not have compact support.
@@
(2.16) $|v_{1}(y)|\leq\frac{C(A,R_{0})}{1+|y|}\textrm{ and }|\nabla^{k}v_{1}(y)|\leq\frac{C(A,\beta,k,R_{0})}{1+|y|^{1+\beta}},~k\in\mathbb{N}^{+},$
@@
First of all, we have
$|\eta_{R_{0}}v_{0}|\leq\frac{C(A)}{1+|y|},~|\nabla(\eta_{R_{0}}v_{0})|\leq\frac{C(A,\beta)}{1+|y|^{1+\beta}}$
due to Lemma 2.1. Hence, to prove (2.16) and (2.17), it suffices to prove (2.17). We only consider large $y$, as the smoothness of $w$ is easy to see using the smoothness of $v_{0}$. We can assume that $|y|\geq 3R_{0}$. Since $v_{0}\cdot\nabla\eta_{R_{0}}$ is smooth and has compact support, we have
@@
$=-\frac{1}{2\pi}\int_{B_{2R_{0}}}\int_{0}^{1}\frac{d}{ds}\frac{y-sz}{|y-sz|^{2}}dsv_{0}\cdot\nabla\eta_{R_{0}}(z)dz$
$=-\frac{1}{2\pi}\int_{B_{2R_{0}}}\int_{0}^{1}(-z)\cdot\nabla\frac{y-sz}{|y-sz|^{2}}dsv_{0}\cdot\nabla\eta_{R_{0}}(z)dz.$
Hence,
@@
where we have used that $|y-sz|\geq|y|-s|z|\geq|y|-|z|\geq\frac{1}{3}|y|$ if $|y|\geq 3R_{0}$ and $|z|\leq 2R_{0}$. This shows that
@@
where $\mathbbm{1}$ is the characteristic function and $B_{R_{0}}^{c}$ is the complement of $B_{R_{0}}$. It is then easy to see that for any $r>2$ we have
@@
Combining the above two inequalities, we obtain (2.18). We conclude the whole proof. ∎
@@
where $v$ is the solution to (2.4). Then $v_{2}$ satisfies
(2.20) $\left\{\begin{aligned} &-\Delta v_{2}-\frac{1}{2}v_{2}-\frac{1}{2}y\cdot\nabla v_{2}+v_{1}\cdot\nabla v_{2}+v_{2}\cdot\nabla v_{1}+v_{2}\cdot\nabla v_{2}+\nabla q_{2}=-v_{1}\cdot\nabla v_{1}-F(v_{0},w),\\ &{\rm div\thinspace}v_{2}=0.\end{aligned}\right$.
where
(2.21) $F(v_{0},w)$ $=-\Delta v_{1}-\frac{1}{2}v_{1}-\frac{1}{2}y\cdot\nabla v_{1}$
@@
To show Theorem 2.1, we first prove a slightly weaker version for $v_{2}$. It says that $v_{2}$ is is $L^{p}$ integrable and hence $v_{\text{re}}$ is also $L^{p}$ integrable for any $1<p<2$.
@@
For any $1<p<2$, we can choose $R_{0}=R_{0}(A,\beta,p)$ in the definition of $v_{1}$ (2.14) large enough, such that $v_{2}=v-v_{1}$, the corresponding solution to (2.20) satisfies
@@
Especially, choosing $p=\frac{3}{2}$ and fix a $R_{0}=R_{0}(A,\beta)$ hence the modified profile $v_{1}$, then there exists a constant $C$ depending only on $A$ and $\beta$ such that
$\int_{\mathbb{R}^{2}}|v_{2}|^{2}+|\nabla v_{2}|^{2}dy\leq C(A,\beta).$
@@
Here and in the following $v_{1,j}$ means the $j$-th component of $v_{1}$, $v_{2,j}$ means the $j$-th component of $v_{2}$ for $j=1,2$. Multiplying the equation of $v_{1}$ (2.21) by $v_{2}$, then integrating both sides, it leads to
$\int_{\mathbb{R}^{2}}\partial_{i}v_{2,j}\partial_{i}v_{1,j}dy-\int_{\mathbb{R}^{2}}\frac{1}{2}v_{1}\cdot v_{2}+\frac{1}{2}y_{i}\partial_{i}v_{1,j}v_{2,j}dy$
$=$ $\int_{\mathbb{R}^{2}}\partial_{i}v_{2,j}\partial_{i}v_{1,j}dy+\int_{\mathbb{R}^{2}}-\frac{1}{2}v_{1}\cdot v_{2}+v_{1}\cdot v_{2}+\frac{1}{2}y_{i}\partial_{i}v_{2,j}v_{1,j}dy=\int_{\mathbb{R}^{2}}F(v_{0},w)\cdot v_{2},$
where integration by parts has been used. This is just
(2.24) $\int_{\mathbb{R}^{2}}\partial_{i}v_{2,j}\partial_{i}v_{1,j}dy+\int_{\mathbb{R}^{2}}\frac{1}{2}v_{1}\cdot v_{2}+\frac{1}{2}(y\cdot\nabla v_{2})\cdot v_{1}dy=\int_{\mathbb{R}^{2}}F(v_{0},w)\cdot v_{2}.$
We note that
@@
For the right-hand side of above, we estimate with the help of Lemma 2.2 as follows:
@@
and it follows from (2.17) that
$\int_{\mathbb{R}^{2}}\left(\frac{1}{2}w+\frac{1}{2}y\cdot\nabla w\right)\cdot v_{1}$
$\leq$ $C(A,R_{0})\int_{\mathbb{R}^{2}}\frac{1}{1+|y|}\left|\frac{1}{2}w+\frac{1}{2}y\cdot\nabla w\right|$
@@
Finally, using Lemma 2.2 again, we have
$\int_{\mathbb{R}^{2}}\Delta w\cdot v_{1}=-\int_{\mathbb{R}^{2}}\nabla w\cdot\nabla v_{1}\leq\int_{\mathbb{R}^{2}}\frac{C(A,\beta,R_{0})}{1+|y|^{4}}=C(A,\beta,R_{0}).$
@@
$\int_{\mathbb{R}^{2}}|\nabla v_{2}|^{2}dy+2\int_{\mathbb{R}^{2}}\partial_{i}v_{2,j}\partial_{i}v_{1,j}dy=-\int_{\mathbb{R}^{2}}F(v_{0},w)v_{1}\leq C(A,\beta,R_{0}),$
@@
$\qquad\qquad=-\int_{\mathbb{R}^{2}}\left(v_{1}\cdot\nabla v_{1}-F(v_{0},w)\right)\cdot|v_{2}|^{p-2}v_{2}.$
@@
Estimates for the first five terms on left-hand side of (2.28). For the first term, we have
@@
where in the last inequality, we have used that $p<2$. In above, $v_{2,i}$ is the $i$-th component of $v_{2}$ and $\partial_{j}v_{2,i}=\frac{\partial v_{2,i}}{\partial x_{j}}$ for $i,j=1,2$. Also, we have
$\int_{\mathbb{R}^{2}}-\frac{1}{2}y\cdot\nabla v_{2}\cdot|v_{2}|^{p-2}v_{2}$ $=\int_{\mathbb{R}^{2}}-\frac{1}{4}y\cdot\nabla|v_{2}|^{2}|v_{2}|^{p-2}$
$=\int_{\mathbb{R}^{2}}-\frac{1}{4}y\cdot\nabla|v_{2}|^{2}(|v_{2}|^{2})^{\frac{p-2}{2}}$
@@
$=\frac{1}{p}\int_{\mathbb{R}^{2}}|v_{2}|^{p}$
And using integration by parts, we have by using ${\rm div\thinspace}v_{1}=0$ that
@@
From the above estimates, we have
@@
Importantly, one note that p1>0p-1>0 and $\frac{1}{p}-\frac{1}{2}>0$ due to $1<p<2$.

### v2

@@ 2. The Leray equations and the Energy estimates
Furthermore, we demonstrate that $u_{\text{re}}(x,1)$ is $L^{p}$-integrable for any $p\in(1,+\infty)$. This implies that $u_{\textrm{re}}$ decays like $\frac{1}{1+|x|^{2}}$ in an average sense. The local regularity of $u_{\text{re}}(x,1)$ and $e^{\Delta}u_{0}$ is not a big issue, indeed they are smooth due to the parabolic smoothing effect, once we have some initial local integrability. However, they decay slowly at spatial infinity, and the key point is to get the global integrability to initiate the subsequent analysis. Before stating the main result in this section, we introduce self-similar variables and the Leray equations.
@@ 2.1. Self-similar variables and the Leray equations
Our analysis will be conducted in self-similar variables; we let
(2.1) $y=\frac{x}{\sqrt{t}}~\textrm{and}~s=\log(t).$
We then define $v(y,s)$ by
@@
(2.3) $\left\{\begin{aligned} &\frac{\partial v}{\partial s}-\Delta v-\frac{1}{2}v-\frac{1}{2}y\cdot\nabla v+v\cdot\nabla v+\nabla q=0,\\ &\textrm{ div }v=0,\end{aligned}\right.\textrm{ in }\mathbb{R}^{2}\times(-\infty,\infty).$
@@
(2.4) $\left\{\begin{aligned} &-\Delta v-\frac{1}{2}v-\frac{1}{2}y\cdot\nabla v+v\cdot\nabla v+\nabla q=0,\\ &\textrm{ div }v=0,\end{aligned}\right.\textrm{ in }\mathbb{R}^{2}.$
@@
$e^{t\Delta}u_{0}(x)=\int_{\mathbb{R}^{2}}\frac{1}{4\pi t}e^{-\frac{|x-z|^{2}}{4t}}u_{0}(z)dz.$
Direct calculation shows that $e^{t\Delta}u_{0}$ inherits the self-similarity of $u_{0}$, i.e.,
$\lambda(e^{\lambda^{2}t\Delta}u_{0})(\lambda x)$ $=\lambda\int_{\mathbb{R}^{2}}\frac{1}{4\pi\lambda^{2}t}e^{-\frac{|\lambda x-z|^{2}}{4\lambda^{2}t}}u_{0}(z)dz$
@@
$=\int_{\mathbb{R}^{2}}\frac{1}{4\pi t}e^{-\frac{|x-z|^{2}}{4t}}u_{0}(z)dz=e^{t\Delta}u_{0}(x),$
@@
$\frac{1}{\sqrt{t}}v_{0}(y)=(e^{t\Delta}u_{0})(x).$
Obviously we have
(2.5) $v_{0}(y)=(e^{\Delta}u_{0})\left(\frac{x}{\sqrt{t}}\right)=(e^{\Delta}u_{0})(y).$
@@
Our primary goal is to get a priori estimates for $v_{\text{re}}(y)$. It is easy to see that $v_{0}$ satisfies
@@
(2.8) $\left\{\begin{aligned} &-\Delta v_{\text{re}}-\frac{1}{2}v_{\text{re}}-\frac{1}{2}y\cdot\nabla v_{\text{re}}+v_{0}\cdot\nabla v_{\text{re}}+v_{\text{re}}\cdot\nabla v_{0}+v_{\text{re}}\cdot\nabla v_{\text{re}}+\nabla q=-v_{0}\cdot\nabla v_{0},\\ &{\rm div\thinspace}v_{\text{re}}=0,\end{aligned}\right.\textrm{ in }\mathbb{R}^{2}.$
We define the following function space for later use.
(2.9) $X=\{{\boldsymbol{v}}:\ {\rm div\thinspace}{\boldsymbol{v}}=0,\ {\boldsymbol{v}}\in C^{1}(\mathbb{R}^{2}),~\|{\boldsymbol{v}}\|_{X}<+\infty\},$
where the norm $\|\cdot\|_{X}$ is
(2.10) $\|{\boldsymbol{v}}\|_{X}=\|(1+|x|^{2}){\boldsymbol{v}}\|_{C(\mathbb{R}^{2})}+\|(1+|x|^{2})\nabla{\boldsymbol{v}}\|_{C(\mathbb{R}^{2})}.$
Indeed, this space $X$ will be the space where we use the Leray-Schauder fixed-point theorem to seek a solution $v_{\text{re}}$ to (2.8). We can rather assume that $v_{\text{re}}\in X$ to facilitate later computations; the key point is to obtain enough a priori estimates of $v_{\text{re}}$ in $X$ in terms of the initial data $u_{0}$ only.
The main objective of this section is to prove the following a priori estimates for $v_{\text{re}}$.
@@
For $0<\beta\leq 1$, given the divergence-free, self-similar initial data $u_{0}(x)\in C^{0,\beta}_{loc}$, and
@@
Assume $u$ is a self-similar solution to (1.1) in the sense of Definition 1.1, and that $v_{\textrm{re}}$ $\in X$ is a corresponding solution to (2.8) in which $v_{0}=e^{\Delta}u_{0}$. Then $v_{\text{re}}$ is smooth and there exists a constant $C=C(A,\beta)$ depending only on $A$ and $\beta$, such that
(2.11) $\int_{\mathbb{R}^{2}}|v_{\text{re}}|^{2}+|\nabla v_{\text{re}}|^{2}dx\leq C(A,\beta).$
@@
$\|\nabla q\|_{L^{p}(\mathbb{R}^{2})}\leq C(A,\beta,p).$
@@ 2.2. Main difficulties, key ideas and the outline for the energy estimates
In this section, we give the main ideas for proving the a priori energy estimates in Theorem 2.1.
@@
yet $\|v_{\text{re}}\|_{L^{2}(\mathbb{R}^{2})}$ cannot be controlled at this stage. This situation differs from the three-dimensional case, where there is an additional term $\frac{1}{4}\int_{\mathbb{R}^{2}}|v_{\text{re}}|^{2}$ that appears in the left-hand side of (2.12) and help to close the energy estimates, provided $\|v_{0}\|_{L^{\infty}}$ is small via a suitable cut-off; see also [8, 9].
@@
as $\nabla v_{0}\sim\frac{C}{1+|y|^{1+\beta}}\in L^{2}(\mathbb{R}^{2})$. This provides the initial estimate for the Dirichlet energy of $v_{\text{re}}$.
@@
The left hand side contains an additional term $\int_{\mathbb{R}^{2}}\left(\frac{1}{p}-\frac{1}{2}\right)|v_{\text{re}}|^{p}$, which enables us to close the estimates for $\|v_{\text{re}}\|_{L^{p}}$ for any $1<p<2$ eventually. The key issue is to control $\int_{\mathbb{R}^{2}}(v_{\text{re}}\cdot\nabla v_{0})\cdot|v_{\text{re}}|^{p-2}v_{\text{re}}$ and $\int_{\mathbb{R}^{2}}\nabla q\cdot|v_{\text{re}}|^{p-2}v_{\text{re}}$ as the last term on the right hand side is more manageable. To achieve this, we introduce a modified profile of $v_{0}$ to ensure it is small near the origin. Finally, once the estimate for $\|v_{\text{re}}\|_{L^{p}}$ is established, the Gagliardo–Nirenberg interpolation inequality combined with (2.14) yields the desired $L^{2}$-boundedness of $v_{\text{re}}$. The higher integrability estimate of $\|v_{\text{re}}\|_{L^{p}}$ when $p>2$ will follow from regularity estimates for the Stokes equation and the bootstrapping argument.
@@ 2.3. Proof of the energy estimates
As preliminaries, we begin with the following decay estimates for the solution to the heat equation. It is more or less standard, and we leave the proof to Appendix A.
@@
Let $u_{0}$ be divergence-free on $\mathbb{R}^{2}$ and self-similar, and let $0\leq\beta\leq 1$, denote
@@
Then $v_{0}=e^{\Delta}u_{0}(y)$ defined in (2.5) is smooth and divergence-free. Furthermore, it holds that
@@
then there exists some universal constant $C$ that
@@
where $w$ is given by
(2.18) $w(y)$ $=-\frac{1}{2\pi}\nabla_{y}\int_{\mathbb{R}^{2}}(\ln|y-z|)v_{0}\cdot\nabla\eta_{R_{0}}(z)dz=-\frac{1}{2\pi}\int_{\mathbb{R}^{2}}\frac{y-z}{|y-z|^{2}}v_{0}\cdot\nabla\eta_{R_{0}}(z)dz.$
It follows that $w$ solves
${\rm div\thinspace}w=-v_{0}\cdot\nabla\eta_{R_{0}},$
and hence ${\rm div\thinspace}v_{1}=0$.
@@
Such a correction term $w$ is obviously not unique. Typically, it can be constructed with compact support in the annulus $B_{2R_{0}}\setminus B_{R_{0}}$ using the Bogovskii formula, given that $v_{0}\cdot\nabla\eta_{R_{0}}$ is smooth and supported in that region (cf. [17, Theorem III.3.3]). The representation in (2.18), however, may not have compact support.
@@
(2.19) $|v_{1}(y)|\leq\frac{C(A,R_{0})}{1+|y|}~\textrm{ and }~|\nabla^{k}v_{1}(y)|\leq\frac{C(A,\beta,k,R_{0})}{1+|y|^{1+\beta}},~k\in\mathbb{N}^{+},$
@@
The smoothness of $v_{1}$ and ${\rm div\thinspace}v_{1}=0$ is clear from the construction. We focus on proving (2.19)-(2.21). First of all, it follows Lemma 2.1 that
$|\eta_{R_{0}}v_{0}|\leq\frac{C(A)}{1+|y|}\quad\text{and}\quad|\nabla(\eta_{R_{0}}v_{0})|\leq\frac{C(A,\beta)}{1+|y|^{1+\beta}}.$
Hence, to prove (2.19) and (2.20), it suffices to prove (2.20). We only consider the case for $|y|\geq 3R_{0}$, as the boundedness and smoothness of $w$ inside $|y|\leq 3R_{0}$ is easy to see using the boundedness and smoothness of $v_{0}$. Since $v_{0}\cdot\nabla\eta_{R_{0}}$ is smooth and has compact support, we have
@@
$=-\frac{1}{2\pi}\int_{B_{2R_{0}}}\left(\int_{0}^{1}\frac{d}{ds}\frac{y-sz}{|y-sz|^{2}}ds\right)v_{0}\cdot\nabla\eta_{R_{0}}(z)dz$
$=-\frac{1}{2\pi}\int_{B_{2R_{0}}}\left(\int_{0}^{1}(-z)\cdot\nabla\frac{y-sz}{|y-sz|^{2}}ds\right)v_{0}\cdot\nabla\eta_{R_{0}}(z)dz.$
Hence, if $|y|\geq 3R_{0}$, one has
@@
where we have used that $|y-sz|\geq|y|-s|z|\geq|y|-|z|\geq\frac{1}{3}|y|$ for $|y|\geq 3R_{0}$ and $|z|\leq 2R_{0}$. This shows that
@@
where $\mathbbm{1}_{\Omega}$ is the characteristic function of $\Omega$ and $B_{R_{0}}^{c}$ is the complement of $B_{R_{0}}$. It is then easy to see that for any $r>2$ we have
@@
Combining the above two inequalities, we obtain (2.21). This finishes the proof of the lemma. ∎
@@
where $v$ is the solution to (2.4). It is easy to find that $v_{2}$ satisfies
(2.23) $\left\{\begin{aligned} &-\Delta v_{2}-\frac{1}{2}v_{2}-\frac{1}{2}y\cdot\nabla v_{2}+v_{1}\cdot\nabla v_{2}+v_{2}\cdot\nabla v_{1}+v_{2}\cdot\nabla v_{2}+\nabla q_{2}=-v_{1}\cdot\nabla v_{1}-F(v_{0},w),\\ &{\rm div\thinspace}v_{2}=0,\end{aligned}\right$.
in $\mathbb{R}^{2}$, where using (2.7), we have
(2.24) $-\Delta v_{1}-\frac{1}{2}v_{1}-\frac{1}{2}y\cdot\nabla v_{1}$ $=F(v_{0},w)$
@@
To show Theorem 2.1, we first prove a slightly weaker version for $v_{2}$. It says that we have $H^{1}$-estimate and $L^{p}$-estimate for $v_{2}$ when $1<p<2$. Hence, we also have $H^{1}$-estimate and $L^{p}$-estimate for $v_{\text{re}}$ when $1<p<2$.
@@
Let the divergence-free, self-similar initial data $u_{0}(x)\in C^{0,\beta}_{loc}$ be given, where $0<\beta\leq 1$, and let $A=\|u_{0}\|_{C^{0,\beta}(S^{1})}<+\infty$. For any $1<p<2$, we can choose $R_{0}=R_{0}(A,\beta,p)$ in the definition of $v_{1}$ (2.17) large enough, such that the following a priori estimates hold,
@@
In particular, choose $p=\frac{3}{2}$ and fix a $R_{0}=R_{0}(A,\beta)$ yields that
$\int_{\mathbb{R}^{2}}|v_{2}|^{2}+|\nabla v_{2}|^{2}dy\leq C(A,\beta),$
where the constant $C(A,\beta)$ depends only on $A$ and $\beta$.
@@
Here and in the following $v_{1,j}$ means the $j$-th component of $v_{1}$, $v_{2,j}$ means the $j$-th component of $v_{2}$ for $j=1,2$. Multiplying Equation (2.24) of $v_{1}$ by $v_{2}$, then integrating both sides, it leads to
(2.27) $\int_{\mathbb{R}^{2}}\partial_{i}v_{2,j}\partial_{i}v_{1,j}dy-\int_{\mathbb{R}^{2}}\frac{1}{2}v_{1}\cdot v_{2}+\frac{1}{2}y_{i}\partial_{i}v_{1,j}v_{2,j}dy$
$=$ $\int_{\mathbb{R}^{2}}\partial_{i}v_{2,j}\partial_{i}v_{1,j}dy+\int_{\mathbb{R}^{2}}-\frac{1}{2}v_{1}\cdot v_{2}+v_{1}\cdot v_{2}+\frac{1}{2}y_{i}\partial_{i}v_{2,j}v_{1,j}dy$
$=$ $\int_{\mathbb{R}^{2}}\partial_{i}v_{2,j}\partial_{i}v_{1,j}dy+\int_{\mathbb{R}^{2}}\frac{1}{2}v_{1}\cdot v_{2}+\frac{1}{2}(y\cdot\nabla v_{2})\cdot v_{1}dy=\int_{\mathbb{R}^{2}}F(v_{0},w)\cdot v_{2},$
where integration by parts has been used. We also note that
@@
For the right-hand side of above, with the help of Lemma 2.2, one has
@@
and
$\int_{\mathbb{R}^{2}}\left(\frac{1}{2}w+\frac{1}{2}y\cdot\nabla w\right)\cdot v_{1}\leq$ $C(A,R_{0})\int_{\mathbb{R}^{2}}\frac{1}{1+|y|}\left|\frac{1}{2}w+\frac{1}{2}y\cdot\nabla w\right|$
@@
Finally, using Lemma 2.2 again yields
$\int_{\mathbb{R}^{2}}\Delta w\cdot v_{1}\leq\int_{\mathbb{R}^{2}}\frac{C(A,\beta,R_{0})}{1+|y|^{5}}=C(A,\beta,R_{0}).$
@@
$\int_{\mathbb{R}^{2}}|\nabla v_{2}|^{2}dy+2\int_{\mathbb{R}^{2}}\partial_{i}v_{2,j}\partial_{i}v_{1,j}dy=-\int_{\mathbb{R}^{2}}F(v_{0},w)v_{1}\leq C(A,\beta,R_{0}).$
@@
$=$ $-\int_{\mathbb{R}^{2}}\left(v_{1}\cdot\nabla v_{1}-F(v_{0},w)\right)\cdot|v_{2}|^{p-2}v_{2}.$
@@
Estimates for the first five terms on the left-hand side of (2.31). The estimates of the first five terms on the left-hand side of (2.31) are quite straightforward using integration by parts. For the first term, we have
@@
where the condition $p<2$ guarantees the second-to-last inequality holds. In above, $v_{2,i}$ is the $i$-th component of $v_{2}$ and $\partial_{j}v_{2,i}=\frac{\partial v_{2,i}}{\partial x_{j}}$ for $i,j=1,2$. Next, we have
$\int_{\mathbb{R}^{2}}-\frac{1}{2}y\cdot\nabla v_{2}\cdot|v_{2}|^{p-2}v_{2}$ $=\int_{\mathbb{R}^{2}}-\frac{1}{4}y\cdot\nabla|v_{2}|^{2}(|v_{2}|^{2})^{\frac{p}{2}-1}$
@@
$=\frac{1}{p}\int_{\mathbb{R}^{2}}|v_{2}|^{p}.$
Using integration by parts and ${\rm div\thinspace}v_{1}=0$ yields that
@@
Combining all the above estimates gives
@@
It is important to note that p1>0p-1>0 and $\frac{1}{p}-\frac{1}{2}>0$ provided $1<p<2$.